Chapter 19, Problem 009 An automobile tire has a volume of 1.58 x 102 m3 and contains air at a gauge pressure (pressure above atmospheric pressure) of 162 kPa when the temperature is 0.00°C. What is the gauge pressure of the air in the tires when its temperature rises to 22.1°C and its volume increases to 1.69 x 102 m3 Assume atmospheric pressure is 1.01 x 105 Pa. Number Units The num The number of significant digits is set to 3; the tolerance is +/-2% Click if you would like to Show Work for this question: Qpen Show Work LINK TO TEXT LINK TO SAMPLE PROBLEM
Answer #1

PV/T= constant

here temperature is in Kelvin and Pressure = gauge pressure + atmospheric pressure

P1 =( PV/T)*(T1/V1)

P1= 263*0.0158*295.25/(273.15*0.0169)

P1=265.78

GAUGE PRESSURE= 265.78-101 =164.78 kPa